Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 169     Entry time: 03 May 2023, 13:01
Author: Jan Gruenert 
Group: PRC 
Subject: Status 

The machine tuning over the night had not yet accomplished fully satisfactory beam conditions mainly at the high photon energies for SA1 and SA2,
therefore tuning continued during WED morning.
Beam was handed over to SCS (SASE3) at 9h with about 2.4mJ at 2.5 keV.
Beam was handed over to FXE (SASE1) at 13h with about 530uJ at 19.5 keV.
Tuning for MID (SASE2) at 23 keV is agreed to continue until 14h, currently at a level of 120 uJ. 

Attachment 1: JG_2023-05-03_um_13.10.33.png  1.327 MB  Uploaded 03 May 2023, 14:10  | Hide | Hide all
JG_2023-05-03_um_13.10.33.png
ELOG V3.1.4-7c3fd00